40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 3 19:02:11 2018, gautam, Update, ASS, X arm ASS recovery 
    Reply  Tue Aug 14 00:50:32 2018, gautam, Update, ASS, X arm ASS still not quite right? POXPOY.pngIMG_7108.JPG
Message ID: 14132     Entry time: Fri Aug 3 19:02:11 2018     Reply to this: 14161
Author: gautam 
Type: Update 
Category: ASS 
Subject: X arm ASS recovery 

[koji, gautam]

After I effected the series resistance change for ETMX, the X arm ASS didn't work (i.e. IR transmission would degrade if the servo was run). Today, we succeeded in recovering a functional ASS servo yes.

So both arms have working dither alignment servos now. But remember that the Y arm ASS gains have been set for locking the Y arm with MC2 as the actuator, not ETMY.

Details:

  • Koji pointed out that the demodulated signals from the ETM dither are only used to center the spot on the ETM, and that we should first run the servo with existing settings with the ETM pitch and yaw spot centering loops disabled.
    • This improved TRX level from ~0.8 to 1.1
  • Next, we tried increasing the LO amplitudes by x5 to account for the reduced actuation of the dither on ETMX
    • We then re-enabled the two loops that were earlier disabled.
    • This resulted in TRX degrading very quickly.
  • So we decided to try going back to the nominal LO gains, and reducing the gain of the two ETM spot centering loops.
    • This did the trick, TRX went from 1.1 --> ~1.23, which is the nominal maximum pre-vent value.
  • The snap file used to recover the correct settings to run the dither alignment servos have been updated, the old one has been backed up with today's datestamp.

We then tried to maximize GTRX using the PZT mirrors, but were only successful in reaching a maximum of 0.41. The value I remember from before the vent was 0.5, and indeed, with the IR alignment not quite optimized before we began this work, I saw GTRX of 0.48. But the IR dither servo signals indicate that the cavity axis may have shifted (spot position on the ITM, which is uncontrolled, seems to have drifred significantly, the Pitch signal doesn't stay on the StripTool scale anymore). So we may have to double check that the transmitted beam isn't falling off the GTRX DC PD.

ELOG V3.1.3-